Assume that in a given region, 40% of young males are alcoholics. Find the probability that in a sample of 10 young males,
1. Only one young male is an alcoholic,
2. At most 2 are alcoholics.
3. More than 4 are alcoholics
1
Expert's answer
2020-12-24T12:57:16-0500

1) "P(X=1)=C_{10}^10.4^1(1-0.4)^{10-1}=0.0403."

2) "P(X\le2)=P(X=0)+P(X=1)+P(X=2)="

"=C_{10}^00.4^00.6^{10}+C_{10}^10.4^10.6^9+C_{10}^20.4^20.6^8=0.1673."

3) "P(x>4)=1-P(X\le3)=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)="

"=1-C_{10}^00.4^00.6^{10}-C_{10}^10.4^10.6^9-C_{10}^20.4^20.6^8-C_{10}^30.4^30.6^7=0.3669."